Portfolio Required Return

Suppose you manage a $5.47 million fund that consists of four stocks with the following investments:

Stock Investment Beta
A $220,000 1.50
B 800,000 -0.50
C 1,500,000 1.25
D 2,950,000 0.75

If the market's required rate of return is 11% and the risk-free rate is 3%, what is the fund's required rate of return? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1
Stock Investment Beta Proportion (P) Rf P*Rf
A      220,000.00 1.5 0.04022 15 0.60
B      800,000.00 -0.5 0.14625 -1 -0.15
C 1,500,000.00 1.25 0.27422 13 3.56
D 2,950,000.00 0.75 0.53931 9 4.85
   5,470,000.00 1.00000 TOTAL 8.88

Fund's required rate of return = 8.88%

Note:

1.Rf = (Rm-Rf)*B + Rf ie A = (11-3)*1.5 + 3 = 15% and similarly change Beta for the rest.

2. Proprtion => A: 220000/5470000 = 0.04022 and similarly for the rest